# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

### v2

@@ 6 Discussion and Conclusions
In this study, we investigated the water volume estimation from seismic data. The field measurements were done for a reservoir for which the water table level is controllable and known. Furthermore, we possessed accurate knowledge of the reservoir’s geometry and physical characteristics. Our approach involved testing the field data against a proposed neural network model, developed through a training phase utilizing synthetic data.
@@
A crucial element of our research was the examination of how closely the estimations from seismic data matched with the actual water volumes and water table levels in the controlled reservoir environment. These correlations are key to verifying the accuracy of our neural network model. Quantitative comparison between the seismic data-derived estimates with the ground truth of the reservoir’s water levels and volumes is fundamental to demonstrating the reliability of our approach. It is especially pertinent when considering the potential application of our methods in varied and uncontrolled real-world scenarios.
@@
Future research should encompass larger test sites with heterogeneous materials. From a methodological perspective, approaches to provide insights into the estimation uncertainty, as well as methods to normalize the source wavelet in time, could be potential research directions. Furthermore, depending on the characteristics of the site under investigation, it may be necessary to extend the current physical models. These extensions could involve accommodating anisotropic materials, addressing issues related to partial saturation, and incorporating more complex noise models.
In this study, synthetic databases were constructed using a single arbitrarily selected source location from field measurements. Future research could explore the inclusion of multiple sources, particularly beneficial for addressing challenges associated with complex geometries and material inhomogeneities.

### v3

@@ 6 Discussion and Conclusions
In this study, we investigated water volume estimation using both field and synthetic seismic data. The field measurements were done for a reservoir for which the water table level is controllable and known. Furthermore, we possessed accurate knowledge of the reservoir’s geometry and physical characteristics. Our approach involved testing the field data against a proposed neural network model, developed through a training phase utilizing synthetic data.
@@
A crucial element of our research was the examination of how closely the estimations from seismic data, based on neural networks, matched with the actual water volumes in the controlled reservoir environment. These correlations are key to verifying the accuracy of our neural network model. Quantitative comparison between the seismic data-derived estimates and the ground truth of the reservoir’s water volumes is fundamental to demonstrating the reliability of our approach. It is especially pertinent when considering the potential application of our methods in varied and uncontrolled real-world scenarios. It is expected that when encountering out-of-training-distribution samples, whether due to deviations in material parameters or geometric inaccuracies, the estimation accuracy may deteriorate. Additionally, certain parameters are likely to have a greater impact on estimation errors than others, making sensitivity to specific variables an important consideration in assessing the robustness of the model.
@@
An important next step in research involves exploring larger test sites. From a methodological perspective, approaches to quantify and characterize the uncertainty in neural network-based estimates, as well as methods to normalize the source wavelet in the time domain, could be valuable directions. Furthermore, depending on the characteristics of the site under investigation, it may be necessary to extend the current physical models. These extensions could involve accommodating inhomogeneous and possibly anisotropic materials, addressing the challenges associated with regions of partial saturation, and incorporating more complex noise models. Additional work could also investigate which parameters have the greatest impact on uncertainty in model outputs. Moreover, expanding the synthetic databases to include multiple source locations, instead of relying on a single, arbitrarily selected source, could help address challenges associated with complex geometries and material inhomogeneities.